細胞生物工学 2 単位 (選択)

Cell Biotechnology  長宗秀明 · 教授/ 環境創生工学専攻生命テクノサイエンスコース生物機能工学講座

友安俊文 · 准教授/ 環境創生工学専攻生命テクノサイエンスコース生物機能工学講座

【授業目的】細胞生物工学技術やその応用例を扱ったディベートによって，細胞

工学の知識の拡充と討論スキルの向上を目指す．【授業概要】細胞生物工学に関する技術や周辺領域の知識を深めるため，細胞を

利用した有用物質の生産や医療測定あるいは治療技術に関する最新の論文を題材として学生間のディベートを行う．また同時に討論やコミュニケーションスキルの訓練も行う．

【授業形式】講義【キーワード】細胞，生物工学，討論【先行科目】 [先行科目] 【関連科目】『生化学特論』 (0.8)，『分子生物工学』 (0.8)，『生物反応工学特論』

(0.8)，『生体高分子化学特論』 (0.5)，『酵素学特論』 (0.5) 【履修要件】学部教育における細胞工学及びその関連科目を理解していること．【履修上の注意】 [注意] 【到達目標】

1. 最新の細胞生物工学技術やその応用例を理解することができる

2. 細胞生物工学関連技術について発表 · 討議できる能力を身につける【授業計画】

1. 細胞を用いた有用物質生産技術 1(抗体の生産)

2. 細胞を用いた有用物質生産技術 2(酵素の生産)

3. 細胞を用いた有用物質生産技術 3(リンホカインの生産)

4. 細胞を用いた有用物質生産技術 4(サイトカインの生産)

5. 細胞を用いた有用物質生産技術 5(生理活性タンパク質の生産)

6. 細胞を用いた検査技術 1(細胞障害性の測定)

7. 細胞を用いた検査技術 2(シグナル伝達経路の解析)

8. 細胞を用いた検査技術 3(トランスクリプトーム解析)

9. 細胞を用いた検査技術 4(ゲノム異常/SNPs 検査)

10. 細胞を用いた検査技術 5(細胞内感染細菌/ウイルスの検査)

11. 細胞を用いた治療技術 1(人工皮膚)

12. 細胞を用いた治療技術 2(免疫細胞療法)

13. 細胞を用いた治療技術 3(細胞移植/臓器移植)

14. 細胞を用いた治療技術 4(クローン発生技術)

15. 細胞を用いた治療技術 5(ES/iPS 細胞)

【成績評価基準】授業最終日に課すレポート 50%，授業中の発表 50%で評価する．【教科書】授業中に紹介する【参考書】授業中に紹介する【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216661 【対象学生】他学科学生も履修可能【連絡先】

⇒ 各々2時間の予習と復習が各2時間の授業の理解と単位取得のために必要である．

Cell Biotechnology 2 units (selection)

Hideaki Nagamune · Professor / Biological Functions, Biological Science and Technology, Earth and Life Environmental Engineering Toshifumi Tomoyasu · Associate Professor / Biological Functions, Biological Science and Technology, Earth and Life Environmental Engineering

Target⟩ This class aims to expand the knowledge on cell biotechnology and improve the skill of debate through the debate dealing with cell biotechnology and its applications.

Outline⟩ Debates among students dealing with the latest reports on technology and application concerned with production of useful materials, medical assay, and medical treatment using various cells are held in order to deepen the knowledge on cell biotechnology and its peripheral field. Training on skills of debate and communication is simultaneously carried out.

Style⟩ Lecture �������⟩ cell, biotechnology, debate Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ “Advanced Biochemistry”(0.8), “Molecular Biotechnology”

(0.8), “Technology for Bioreaction”(0.8), “Biological macromolecular chemistry” (0.5), “Advanced enzyme engineering”(0.5)

Requirement⟩ Students are required to have a good understanding of undergraduate-

level of cell technology and related subjects Notice⟩ [注意]

Goal⟩

1. To understand the latest cell biotechnology and its application examples

2. To gain the ability of debate on technologies concerned in cell biotechnology

Schedule⟩

1. Cell biotechnology on production of useful materials 1(Production of antibodies)

2. Cell biotechnology on production of useful materials 2(Production of enzymes)

3. Cell biotechnology on production of useful materials 3(Production of lymphokines)

4. Cell biotechnology on production of useful materials 4(Production of cytokines)

5. Cell biotechnology on production of useful materials 5(Production of bioactive proteins)

6. Cell biotechnology on medical or industrial assay 1(Measurement of cytotoxicity)

7. Cell biotechnology on medical or industrial assay 2(Analysis of signal transduction pathways)

8. Cell biotechnology on medical or industrial assay 3(Analysis of transcriptosome)

9. Cell biotechnology on medical or industrial assay 4(Diagnostics of genome disorder/SNPs)

10. Cell biotechnology on medical or industrial assay 5(Diagnostics of intracellular bacteria/viruses)

11. Cell biotechnology on medical treatment 1(Artificial skin)

12. Cell biotechnology on medical treatment 2(Immune cell therapy)

13. Cell biotechnology on medical treatment 3(Cell transplant/Organ transplant)

14. Cell biotechnology on medical treatment 4(Techniques for clone development)

15. Cell biotechnology on medical treatment 5(ES cells/iPS cells) Evaluation Criteria⟩ Assignments count 50%,Presentation count 50% Textbook⟩ To be introduced in the class Reference⟩ To be introduced in the class

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216661

Student⟩ Able to be taken by student of other department

Contact⟩

⇒ Preparation and review of two hours each are required for understanding of every class with two hours and taking the credits.